OpenStudy (anonymous):

Calculate the follow INTEGRAL:

OpenStudy (anonymous):

\[\int\limits_{}^{}(tg^3(x)+tg(x))dx\]

OpenStudy (anonymous):

Are you given anything about g(x)? use integration by parts.

OpenStudy (anonymous):

what in the world is tg? is it tangent?

OpenStudy (anonymous):

@eliassaab @thomaster @texaschic101 @undeadknight26 @phi @PeterPan @kropot72 @Luigi0210 @Lena772 @mathslover @NaomiBell1997

OpenStudy (anonymous):

tg is tan "In my country"

OpenStudy (anonymous):

\[=\int\limits_{}^{}tg^3(x)+\int\limits_{}^{}tg(x)=\int\limits_{}^{}tg^2(x)*tg(x)+\int\limits_{}^{}tg(x)\] \[=\int\limits_{}^{}(\sec^2(x)-1)tg(x)+\int\limits_{}^{}tg(x)=\int\limits_{}^{}\sec^2(x)tg(x)\] Now as long as (tg(x))'=sec^2(x) we substitute tg(x) with u: \[\int\limits_{}^{}(u)du=\frac{ tg^2(x) }{ 2 }+C\] THANKS VERY MUCH GUYS FOR YOUR SUPPORT (SARCASM)

OpenStudy (anonymous):

If just somebody can find a "simple" way to do it ?

OpenStudy (anonymous):

to be fair, it did take you quite a while to respond to my minor query. oh well, as long as you already have it done :)

OpenStudy (anonymous):

and I don't get it, how simple is simple? Like you wouldn't need to use substitution?

OpenStudy (anonymous):

It was a hard substitution to do

OpenStudy (anonymous):

also, no need to split the integral into two. just factor out one of the tg(x) \[\large\int [\tan^3(x) + \tan(x)]dx = \int \tan(x) [\tan^2(x)+1]dx = \int \tan(x)\sec^2(x)\] heck, even a substitution of u = sec(x) would work XD

OpenStudy (anonymous):

@PeterPan You're right